OpenStudy (anonymous):

Given f(x)= -2 and g(x) = 4x-8, find h(x)= f(x) +g(x)

jimthompson5910 (jim_thompson5910):

h(x)= f(x) +g(x) h(x)= -2 + 4x - 8 Now combine like terms

OpenStudy (anonymous):

Thanks so much!

jimthompson5910 (jim_thompson5910):

np